# Ambiguous Online Learning

### v2

@@ 1 Introduction
The condition $y\in\alpha$ requires that the prediction is sound, whereas the condition $\alpha\subseteq h^{*}(x)$ requires that the prediction is “complete”, i.e. no weaker than what knowledge of $h^{*}$ would allow [^4] If we interpret a prediction $\alpha$ as making the logical statement $y\in\alpha$ about the following label $y$, then a prediction is sound iff the statement is true and complete iff the statement is as strong as can be reasonably expected.. Any prediction for which the success criterion doesn’t hold is considered to be a “mistake”. In classical online learning theory, the maximin number of mistakes for any hypothesis class, assuming realizability, is given by the Littlestone dimension. In particular, the asymptotic as a function of the time horizon $N$ can be either $O(1)$ (when the Littlestone dimension is finite) or $\Omega(N)$ (when the Littlestone dimension is infinite). On the other hand, for ambiguous online learning, we prove a trichotomy of possible mistake bounds: depending on the hypothesis class, the number of mistakes can be $O(1)$, $\tilde{\Theta}(\sqrt{N})$ or $\Omega(N)$. In the $O(1)$ case, the exact number of mistakes is given by a combinatorial invariant that we call “ambiguous Littlestone dimension”. In the $\tilde{\Theta}(\sqrt{N})$ case, the coefficient can be bounded using the ordinary Littlestone dimension for the label set $2^{Y}$, and another combinatorial “dimension” associated with the set of subsets of $Y$ that hypotheses can produce.
@@
Apple tasting (see Helmbold et al. (2000), Raman et al. (2024), Chase and Mehalel (2024)) is a variant of online learning with labels $Y=\{0,1\}$ where the learner only receives feedback when predicting $1$. This setting is known to have a trichotomy of mistake bounds similar to our own. In Section 7 we show that there is a formal relationship between the two frameworks. Specifically, there is a way to translate an apple tasting hypothesis class into an ambiguous online learning hypothesis class in a way that preserves the minimax mistake bound, explaining why the trichotomies are the same.
@@ 2 Setting
Formally, a (deterministic) learner is a mapping $A:(X\times Y)^{*}\times X\to 2^{Y}$. Here, the notation $\Sigma^{*}$ means words over the alphabet $\Sigma$. In this case, the first argument of $A$ stands for the past observed history of instances and labels, and the second argument for the current instance for which we need to predict the label.
Given a learner $A$, a hypothesis $h:X\to 2^{Y}$, a number [^8] We use the convention $0\in\mathbb{N}$. $n\in\mathbb{N}$ and a trace $xy\in(X\times Y)^{n}$, we define the set of mistakes that $A$ makes on $xy$ relative to $h$ by
@@
Here, $w_{(:k)}$ stands for the prefix of length $k$ of a word or sequence $w$, $w_{k}$ stands for the $k$-th symbol in the word (starting from 0, i.e. $w=w_{0}w_{1}\dots$), and $x$ and $y$ are the projections of $xy$ to $X^{n}$ and $Y^{n}$ respectively.
@@
Given $h:X\to 2^{Y}$ and $\in(X\times Y)^{n}, xyxy$ is said to be compatible with $h$ when for all $k<n$, $y_{k}\in h(x_{k})$. We denote by $\mathcal{C}_{h}(n)$ the set of all such $xy$. We will also use the notations
@@
It’s also possible to consider a randomized learner $R:(X\times Y)^{*}\times X\to\Delta 2^{Y}$. Here, $\Delta\Sigma$ is our notation for the set of probability distribution over a set $\Sigma$. In this case, we can analogously define the expected number of mistakes. Given a randomized learner $R$, a hypothesis $h:X\to 2^{Y}$, some $n\in\mathbb{N}$ and a trace $xy\in(X\times Y)^{n}$, we define the expected number of mistakes that $R$ makes on $xy$ relative to $h$ to be [^9] Instead of static traces $xy$, it is possible to consider a setting in which the adversary can dynamically choose instances depending on the learner’s predictions. Such a setting can only be harder for a randomized learner, but is equivalent for a deterministic learner. Therefore, the maximin mistake bound in that setting would lie between the lower bound of Proposition 4.7 and the upper bound of Theorem 4.1.
@@
We also define $\mathcal{EM}_{h}^{R}(N)$, $\mathcal{EM}_{\mathcal{H}}^{R}(N)$ and $\mathcal{EM}_{\mathcal{H}}^{*}(N)$ analogously to the deterministic case. Obviously, $\mathcal{EM}^{*}_{\mathcal{H}}(N)\leq\mathcal{M}^{*}_{\mathcal{H}}(N)$, but this inequality might be strict. In this work, we mostly focus on deterministic learners, but also demonstrate a lower bound on $\mathcal{EM}^{*}_{\mathcal{H}}(N)$ which limits how much it can improve on $\mathcal{M}^{*}_{\mathcal{H}}(N)$ to a factor that is constant in $N$.
@@ 3 Combinatorial Invariants
In any $v\in V$ there are at most n1n-1 0s, and between the 0s we have stretches of $1$s of length at most n1n-1. Hence, $|v|\leq n-1+n(n-1)=n^{2}-1$ and $\operatorname{dep}(T)\leq n^{2}-1$.
@@
Given $A\subseteq Y$, define the $\Lambda$-hull of $A$ to be
@@
Given $A\subseteq Y$, define the $\Lambda$-complexity of $A$ by
@@
In words, the $\Lambda$-hull of $A$ is the minimal element of $\Lambda$ containing $A$. The $\Lambda$-complexity of $A$ is the minimal size of a subset $B$ of $A$ s.t. the $\Lambda$-hull of $B$ contains all of $A$. The lower pivot dimension is the maximal $\Lambda$-complexity of any set in $\Lambda$, whereas the upper pivot dimension is the maximal $\Lambda$-complexity of any arbitrary subset of $Y$.
@@ 5 Ambiguous Optimal Algorithm
The $w$-weighted ambiguous Littlestone dimension of $\mathcal{H}$ (denoted $\mathrm{AL}_{w}(\mathcal{H})$) is defined to be the maximal $w$-weighted rank of an ambiguous shattered $\mathcal{H}$-tree, or $\infty$ if the rank is unbounded. [^15] Curiously, Filmus et al. (2023) also have a notion of Littlestone dimension weighted by a mapping $\mathcal{H}\to\mathbb{N}$. However, in their setting the weights represent departure from realizability rather than extra mistakes. We also denote by $\mathrm{AL}_{w}(\mathcal{H},n)$ the maximal rank of an ambiguous shattered $\mathcal{H}$-tree of depth at most $n$. With these ingredients, we can describe the Ambiguous Optimal Algorithm (Algorithm 1), which is an extension of Littlestone’s Standard Optimal Algorithm. The algorithm explicitly depends on the parameter $N$.
@@
Here and everywhere, $\mathbf{1}_{\phi}$ stands for $1$ when $\phi$ is true and 0 when $\phi$ is false. We then choose a prediction by minimaxing the weighted ambiguous Littlestone dimension of the remaining hypotheses on the next step. That is:
@@ 6 Weighted Aggregation Algorithm
We start from describing an algorithm for finite hypothesis classes (i.e., assuming $|\mathcal{H}|<\infty$). This will later be modified to produce the algorithm for Theorem 4.4. We call it the Weighted Aggregation Algorithm (Algorithm 2), denoted by $A_{\text{WAA}}$.
@@
Now suppose that $\mathcal{H}$ might be infinite, but $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})<\infty$. Fix the time horizon $N\in\mathbb{N}$. We then construct a new hypothesis class $\mathcal{H}^{((N))}$ with domain $X^{((N))}:=X^{<N}\times X$. Here, the notation $\Sigma^{<N}$ (resp. $\Sigma^{\leq N}$) means words over alphabet $\Sigma$ of length less than $N$ (resp. less than or equal to $N$).
@@
Consider any $S\subseteq[N]$ with $|S|\leq\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and $f:S\to Y^{\mathfrak{P}}$. Define $h_{f}:X^{((N))}\to\Lambda(\mathcal{H})$ and $h^{*}_{f}:X^{\leq N}\to(X\times Y^{\mathfrak{P}})^{*}$ by simultaneous recursion via
@@
$\mathcal{H}^{((N))}:=\{h_{f}\mid S\subseteq[N]\text{ with }|S|\leq\mathrm{L}_{\mathfrak{P}}(\mathcal{H})\text{ and }f:S\to Y^{\mathfrak{P}}\}$
The hypothesis class $\mathcal{H}^{((N))}$ can “imitate” the original class $\mathcal{H}$ in the following sense.
@@
For any $x\in X^{N}$ and $h\in\mathcal{H}$, if for all $n<N$ it holds that $h(x_{n})\neq\emptyset$, then there exists $h^{\prime}\in\mathcal{H}^{((N))}$ s.t. for all $n<N$, $h^{\prime}(x_{:n+1})=h(x_{n})$.
(See Appendix F for the proof.) Now consider the learner $A^{\prime}_{\text{WAA}}$ which is the WAA associated with $\mathcal{H}^{((N))}$. Using Theorem 6.1 and Lemma 6.2, it is straightforward to see that $A^{\prime}_{\text{WAA}}$ satisfies the upper bound of Theorem 4.4 w.r.t the hypothesis class $\mathcal{H}$.
@@ 7 Apple Tasting and Additional Lower Bounds
There is a close connection between ambiguous online learning and so-called apple tasting (see Helmbold et al. (2000)). In apple tasting, there is a hypothesis class $\mathcal{H}\subseteq\{X\to\{0,1\}\}$ and the learner’s output is also in $\{0,1\}$. As in classical online learning, whenever the learner’s prediction differs from the true label, this counts as a mistake. In contrast to classical online learning, when the learner’s prediction is 0, the learner receives no feedback (whereas if the learner’s prediction is $1$, the true label is revealed). We will use the notation $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$ for apple tasting learners, where “no feedback” is represented by $1$.
@@ 8 Conclusion
Qualitatively, instead of the two possible behaviors of classical theory ($O(1)$ mistakes and $N$ mistakes), we have three possible behaviors, with the new option of $O(\sqrt{N})$ mistakes. The upper bound of this new option is attained using the WAA, an algorithm reminiscent of weighted majority. The coefficient depends both on the classical Littlestone dimension $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and the combinatorial invariant $\underline{\mathrm{PD}}(\mathcal{H})$. As in classical theory, there is at most a constant (w.r.t. $N$) gap between the mistake bounds of deterministic and randomized learners. However, our current best characterization of this gap involves yet another new combinatorial invariant, namely $\overline{\mathrm{PD}}(\mathcal{H})$. There are many natural questions for further study:
@@ Appendix A Properties of Combinatorial Invariants
To prove Proposition 3.4, we traverse a path going down from the root. At each vertex $v$ on the path, there will be a leaf label $h$ for which the $E_{0}$-edge of $v$ is relevant, and we’ll continue the path through one of the other edges which possesses an edge label outside of $h(\mathbf{x}_{v})$. The leaf labels we encounter in this way are all different.
@@
In order to calculate $\underline{\mathrm{PD}}(\Lambda)$ in Example 3.9, we observe that every $A\in\Lambda$ is a hyperrectangle and the $\Lambda$-hull of its two extreme corners equals $A$. In order to calculate $\overline{\mathrm{PD}}(\Lambda)$, for any $A\subseteq[n]^{d}$ we take, for each dimension, a point in $A$ with minimal coordinate along this dimension and a point in $A$ with maximal coordinate. These are $2d$ points whose $\Lambda$-hull is $A$.
@@ Appendix C Mistakes and the Ambiguous Littlestone Dimension
We conclude that $\mathcal{M}_{\mathcal{H}}^{A}(n)\geq\mathcal{M}_{h}^{A}(n)\geq\mathcal{M}_{h}^{A}(m)\geq|R_{T}(v_{m})|\geq r$. Since this holds for any $A$, we get $\mathcal{M}_{\mathcal{H}}^{*}(n)\geq r$. ∎
@@
Here, $A$ is some learner and $w^{\overline{xy}}$ is a function from $\mathcal{H}_{\text{uf}}^{\overline{xy}}$ to $\mathbb{N}$.
@@ Appendix D Trimming Shattered Trees
The idea of the proof is judiciously selecting $N$ edges out of the $E_{0}$-path starting at the root, and then using induction to trim the subtrees attached to these edges.
@@
Now, assume the claim holds for some $N$. Consider an $\mathcal{H}$-tree $T=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ s.t. $\operatorname{rank}(T)\geq(N+1)^{2}$. Construct the sequence $\{v_{i}\in V\}_{i\leq d}$, for some $d\in\mathbb{N}$, as follows:

### v3

@@ 1 Introduction
The condition $y\in\alpha$ requires that the prediction is sound, whereas the condition $\alpha\subseteq h^{*}(x)$ requires that the prediction is “complete”, i.e. no weaker than what knowledge of $h^{*}$ would allow [^4] If we interpret a prediction $\alpha$ as making the logical statement $y\in\alpha$ about the following label $y$, then a prediction is sound iff the statement is true and complete iff the statement is as strong as can be reasonably expected.. Any prediction for which the success criterion doesn’t hold is considered to be a “mistake”. In classical online learning theory, the maximin number of mistakes for any hypothesis class, assuming realizability, is given by the Littlestone dimension. In particular, the asymptotic as a function of the time horizon NN can be either $O(1)$ (when the Littlestone dimension is finite) or $\Omega(N)$ (when the Littlestone dimension is infinite). On the other hand, for ambiguous online learning, we prove a trichotomy of possible mistake bounds: depending on the hypothesis class, the number of mistakes can be $O(1)$, $\tilde{\Theta}(\sqrt{N})$ or $\Omega(N)$. In the $O(1)$ case, the exact number of mistakes is given by a combinatorial invariant that we call “ambiguous Littlestone dimension”. In the $\tilde{\Theta}(\sqrt{N})$ case, the coefficient can be bounded using the ordinary Littlestone dimension for the label set $2^{Y}$, and another combinatorial “dimension” associated with the set of subsets of $Y$ that hypotheses can produce.
@@
Apple tasting (see Helmbold et al. (2000), Raman et al. (2024), Chase and Mehalel (2024)) is a variant of online learning with labels $Y=\{0,1\}$ where the learner only receives feedback when predicting 11. This setting is known to have a trichotomy of mistake bounds similar to our own. In Section 7 we show that there is a formal relationship between the two frameworks. Specifically, there is a way to translate an apple tasting hypothesis class into an ambiguous online learning hypothesis class in a way that preserves the minimax mistake bound, explaining why the trichotomies are the same.
@@ 2 Setting
Formally, a (deterministic) learner is a mapping $A:(X\times Y)^{*}\times X\to 2^{Y}$. Here, the notation $\Sigma^{*}$ means words over the alphabet $\Sigma$. In this case, the first argument of AA stands for the past observed history of instances and labels, and the second argument for the current instance for which we need to predict the label.
Given a learner AA, a hypothesis $h:X\to 2^{Y}$, a number [^8] We use the convention $0\in\mathbb{N}$. $n\in\mathbb{N}$ and a trace $xy\in(X\times Y)^{n}$, we define the set of mistakes that AA makes on xyxy relative to $h$ by
@@
Here, $w_{(:k)}$ stands for the prefix of length $k$ of a word or sequence $w$, $w_{k}$ stands for the $k$-th symbol in the word (starting from 0, i.e. $w=w_{0}w_{1}\dots$), and $x$ and $y$ are the projections of xyxy to $X^{n}$ and $Y^{n}$ respectively.
@@
Given $h:X\to 2^{Y}$ and $\in(X\times Y)^{n}, xyxy$ is said to be compatible with $h$ when for all $k<n$, $y_{k}\in h(x_{k})$. We denote by $\mathcal{C}_{h}(n)$ the set of all such xyxy. We will also use the notations
@@
It’s also possible to consider a randomized learner $R:(X\times Y)^{*}\times X\to\Delta 2^{Y}$. Here, $\Delta\Sigma$ is our notation for the set of probability distribution over a set $\Sigma$. In this case, we can analogously define the expected number of mistakes. Given a randomized learner $R$, a hypothesis $h:X\to 2^{Y}$, some $n\in\mathbb{N}$ and a trace $xy\in(X\times Y)^{n}$, we define the expected number of mistakes that $R$ makes on xyxy relative to $h$ to be [^9] Instead of static traces xyxy, it is possible to consider a setting in which the adversary can dynamically choose instances depending on the learner’s predictions. Such a setting can only be harder for a randomized learner, but is equivalent for a deterministic learner. Therefore, the maximin mistake bound in that setting would lie between the lower bound of Proposition 4.7 and the upper bound of Theorem 4.1.
@@
We also define $\mathcal{EM}_{h}^{R}(N)$, $\mathcal{EM}_{\mathcal{H}}^{R}(N)$ and $\mathcal{EM}_{\mathcal{H}}^{*}(N)$ analogously to the deterministic case. Obviously, $\mathcal{EM}^{*}_{\mathcal{H}}(N)\leq\mathcal{M}^{*}_{\mathcal{H}}(N)$, but this inequality might be strict. In this work, we mostly focus on deterministic learners, but also demonstrate a lower bound on $\mathcal{EM}^{*}_{\mathcal{H}}(N)$ which limits how much it can improve on $\mathcal{M}^{*}_{\mathcal{H}}(N)$ to a factor that is constant in NN.
@@ 3 Combinatorial Invariants
In any $v\in V$ there are at most n1n-1 0s, and between the 0s we have stretches of 11s of length at most n1n-1. Hence, $|v|\leq n-1+n(n-1)=n^{2}-1$ and $\operatorname{dep}(T)\leq n^{2}-1$.
@@
Given $A\subseteq Y$, define the $\Lambda$-hull of AA to be
@@
Given $A\subseteq Y$, define the $\Lambda$-complexity of AA by
@@
In words, the $\Lambda$-hull of AA is the minimal element of $\Lambda$ containing AA. The $\Lambda$-complexity of AA is the minimal size of a subset $B$ of AA s.t. the $\Lambda$-hull of $B$ contains all of AA. The lower pivot dimension is the maximal $\Lambda$-complexity of any set in $\Lambda$, whereas the upper pivot dimension is the maximal $\Lambda$-complexity of any arbitrary subset of $Y$.
@@ 5 Ambiguous Optimal Algorithm
The $w$-weighted ambiguous Littlestone dimension of $\mathcal{H}$ (denoted $\mathrm{AL}_{w}(\mathcal{H})$) is defined to be the maximal $w$-weighted rank of an ambiguous shattered $\mathcal{H}$-tree, or $\infty$ if the rank is unbounded. [^15] Curiously, Filmus et al. (2023) also have a notion of Littlestone dimension weighted by a mapping $\mathcal{H}\to\mathbb{N}$. However, in their setting the weights represent departure from realizability rather than extra mistakes. We also denote by $\mathrm{AL}_{w}(\mathcal{H},n)$ the maximal rank of an ambiguous shattered $\mathcal{H}$-tree of depth at most $n$. With these ingredients, we can describe the Ambiguous Optimal Algorithm (Algorithm 1), which is an extension of Littlestone’s Standard Optimal Algorithm. The algorithm explicitly depends on the parameter NN.
@@
Here and everywhere, $\mathbf{1}_{\phi}$ stands for 11 when $\phi$ is true and 0 when $\phi$ is false. We then choose a prediction by minimaxing the weighted ambiguous Littlestone dimension of the remaining hypotheses on the next step. That is:
@@ 6 Weighted Aggregation Algorithm
We start from describing an algorithm for finite hypothesis classes (i.e., assuming $|\mathcal{H}|<\infty$). This will later be modified to produce the algorithm for Theorem $4.$ We call it the Weighted Aggregation Algorithm (Algorithm 2), denoted by $A_{\text{WAA}}$.
@@
Now suppose that $\mathcal{H}$ might be infinite, but $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})<\infty$. Fix the time horizon $N\in\mathbb{N}$. We then construct a new hypothesis class $\mathcal{H}^{(N)}$ with domain $X^{(N)}:=X^{<N}\times X$. Here, the notation $\Sigma^{<N}$ (resp. $\Sigma^{\leq N}$) means words over alphabet $\Sigma$ of length less than NN (resp. less than or equal to NN).
@@
Consider any $S\subseteq[N]$ with $|S|\leq\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and $f:S\to Y^{\mathfrak{P}}$. Define $h_{f}:X^{(N)}\to\Lambda(\mathcal{H})$ and $h^{*}_{f}:X^{\leq N}\to(X\times Y^{\mathfrak{P}})^{*}$ by simultaneous recursion via
@@
$\mathcal{H}^{(N)}:=\{h_{f}\mid S\subseteq[N]\text{ with }|S|\leq\mathrm{L}_{\mathfrak{P}}(\mathcal{H})\text{ and }f:S\to Y^{\mathfrak{P}}\}$
The hypothesis class $\mathcal{H}^{(N)}$ can “imitate” the original class $\mathcal{H}$ in the following sense.
@@
For any $x\in X^{N}$ and $h\in\mathcal{H}$, if for all $n<N$ it holds that $h(x_{n})\neq\emptyset$, then there exists $h^{\prime}\in\mathcal{H}^{(N)}$ s.t. for all $n<N$, $h^{\prime}(x_{:n+1})=h(x_{n})$.
(See Appendix F for the proof.) Now consider the learner $A^{\prime}_{\text{WAA}}$ which is the WAA associated with $\mathcal{H}^{(N)}$. Using Theorem 6.1 and Lemma 6.2, it is straightforward to see that $A^{\prime}_{\text{WAA}}$ satisfies the upper bound of Theorem 4.4 w.r.t the hypothesis class $\mathcal{H}$.
@@ 7 Apple Tasting and Additional Lower Bounds
There is a close connection between ambiguous online learning and so-called apple tasting (see Helmbold et al. (2000)). In apple tasting, there is a hypothesis class $\mathcal{H}\subseteq\{X\to\{0,1\}\}$ and the learner’s output is also in $\{0,1\}$. As in classical online learning, whenever the learner’s prediction differs from the true label, this counts as a mistake. In contrast to classical online learning, when the learner’s prediction is 0, the learner receives no feedback (whereas if the learner’s prediction is 11, the true label is revealed). We will use the notation $A:(X\times\{0,1\})^{*}\times X\to\{0,1\}$ for apple tasting learners, where “no feedback” is represented by 11.
@@ 8 Conclusion
Qualitatively, instead of the two possible behaviors of classical theory ($O(1)$ mistakes and NN mistakes), we have three possible behaviors, with the new option of $O(\sqrt{N})$ mistakes. The upper bound of this new option is attained using the WAA, an algorithm reminiscent of weighted majority. The coefficient depends both on the classical Littlestone dimension $\mathrm{L}_{\mathfrak{P}}(\mathcal{H})$ and the combinatorial invariant $\underline{\mathrm{PD}}(\mathcal{H})$. As in classical theory, there is at most a constant (w.r.t. NN) gap between the mistake bounds of deterministic and randomized learners. However, our current best characterization of this gap involves yet another new combinatorial invariant, namely $\overline{\mathrm{PD}}(\mathcal{H})$. There are many natural questions for further study:
@@ Appendix A Properties of Combinatorial Invariants
To prove Proposition 3.4, we traverse a path going down from the root. At each vertex $v$ on the path, there will be a leaf label $h$ for which the $E_{0}$-edge of $v$ is relevant, and we’$l$ continue the path through one of the other edges which possesses an edge label outside of $h(\mathbf{x}_{v})$. The leaf labels we encounter in this way are all different.
@@
In order to calculate $\underline{\mathrm{PD}}(\Lambda)$ in Example 3.9, we observe that every $A\in\Lambda$ is a hyperrectangle and the $\Lambda$-hull of its two extreme corners equals AA. In order to calculate $\overline{\mathrm{PD}}(\Lambda)$, for any $A\subseteq[n]^{d}$ we take, for each dimension, a point in AA with minimal coordinate along this dimension and a point in AA with maximal coordinate. These are $2d$ points whose $\Lambda$-hull is AA.
@@ Appendix C Mistakes and the Ambiguous Littlestone Dimension
We conclude that $\mathcal{M}_{\mathcal{H}}^{A}(n)\geq\mathcal{M}_{h}^{A}(n)\geq\mathcal{M}_{h}^{A}(m)\geq|R_{T}(v_{m})|\geq r$. Since this holds for any AA, we get $\mathcal{M}_{\mathcal{H}}^{*}(n)\geq r$. ∎
@@
Here, AA is some learner and $w^{\overline{xy}}$ is a function from $\mathcal{H}_{\text{uf}}^{\overline{xy}}$ to $\mathbb{N}$.
@@ Appendix D Trimming Shattered Trees
The idea of the proof is judiciously selecting NN edges out of the $E_{0}$-path starting at the root, and then using induction to trim the subtrees attached to these edges.
@@
Now, assume the claim holds for some NN. Consider an $\mathcal{H}$-tree $T=(V,v_{0},P,\mathbf{x},\mathbf{y},\mathbf{h},E_{0})$ s.t. $\operatorname{rank}(T)\geq(N+1)^{2}$. Construct the sequence $\{v_{i}\in V\}_{i\leq d}$, for some $d\in\mathbb{N}$, as follows: